# thanks: These authors contributed equallythanks: These authors contributed equally

### v2

@@ Appendix A Appendix: PXP Model
Appendix S3 A4. PXP model
The scar subspace of the PXP model can also be approximately described using the Hamiltonian decomposition in Eq. (S1) and the Liouvillean in Eq. (S4). Within this approximation, we can directly apply our analysis above by simply replacing the expression for the self-energy:
$\hat{\Gamma}_{\mathrm{PXP}}=\sum_{p_{\mathrm{S}}}|p_{\mathrm{S}}\rangle\langle p_{\mathrm{S}}|\sqrt{\prod_{j=1}^{N-1}\left\langle p_{\mathrm{S}}\right|\hat{P}_{j}\hat{P}_{j+1}\left|p_{\mathrm{S}}\right\rangle}=\sqrt{J\prod_{j=1}^{N-1}\hat{P}_{j}\hat{P}_{j+1}},$ (S24)
where $\hat{P}_{j}=(1-\hat{\sigma}^{z}_{j})/2$. This approximation successfully explains some features of the PXP Liouvillean spectra, e.g., the asymmetry about the imaginary axis, which can be traced back to the following property of the self-energy:
$|\langle E_{n_{\mathrm{S}}}|\hat{\Sigma}_{\mathrm{PXP}}|E_{n_{\mathrm{S}}}\rangle|$ $\neq|\langle E_{N-n_{\mathrm{S}}}|\hat{\Sigma}_{\mathrm{PXP}}|E_{N-n_{\mathrm{S}}}\rangle|.$ (S25)
By contrast, the analogous Eq. (S25) in the SCL model is obeyed as an equality, consistent with the symmetry of the SCL Liouvillean spectrum.
However, the above approximation is not expected to be quantitatively accurate as it ignores the effect of the Rydberg constraint on the PXP Hilbert space. The constraint makes it challenging to analytically perform the perturbation theory for the PXP Liouvillean. Furthermore, to obtain quantitatively accurate results, the QMBS Hamiltonian $\hat{S}_{0}$ should be replaced by an effective spin-1 paramagnet in the dimerized picture [24].
Appendix S4 B. Symmetries of Liouvillean eigenmodes
@@
In this section, we study the symmetry properties of the Liouvillean eigenmodes. Recall that the unperturbed $\mathcal{S}_{0}$ has eigenmodes $\Big{|}\lambda_{(l,s)}^{(0)}\Big{\rangle}\!\!\Big{\rangle}=|E_{s}\rangle\otimes|E_{l+s}\rangle$. Take $N=3$ case as an example. There are $N+1=4$ eigenstates of $\hat{S}_{0}$ in the maximal spin representation:
$|E_{0}\rangle=|---\rangle,\quad\quad|E_{1}\rangle=\frac{1}{\sqrt{3}}\Big{(}|+--\rangle+|-+-\rangle+|--+\rangle\Big{)},$ (S26)
$|E_{2}\rangle=\frac{1}{\sqrt{3}}\Big{(}|++-\rangle+|+-+\rangle+|-++\rangle\Big{)},\quad\quad|E_{3}\rangle=|+++\rangle,$
@@
with energies $-3J$, $-J$, $J$ and $3J$, respectively. The Liouvillean spectrum of $(\mathcal{S}_{0}$ has a seven-layer structure with (imaginary) eigenvalues 00, ±$\pm 2J$, ±$\pm 4J$, ±$\pm 6J$.
@@
The Liouvillean spectrum of $(\mathcal{S}_{0}+\gamma\mathcal{D}^{\prime})$ maintains the seven-layer structure, but now includes eigenvalues with non-zero real parts. In the layer of $l=1$, there are two types of Liouvillean eigenmodes. Both of them are the linear superpositions of $|E_{0}\rangle\otimes|E_{1}\rangle$, $|E_{1}\rangle\otimes|E_{2}\rangle$, and $|E_{2}\rangle\otimes|E_{3}\rangle$, but with different coefficients. One of them is purely imaginary,
$\frac{1}{\sqrt{2}}|E_{0}\rangle\otimes|E_{1}\rangle-\frac{1}{\sqrt{2}}|E_{2}\rangle\otimes|E_{3}\rangle,$ (S27)
@@
$\frac{1}{2}|E_{0}\rangle\otimes|E_{1}\rangle+\frac{1}{2}|E_{2}\rangle\otimes|E_{3}\rangle\pm\frac{1}{\sqrt{2}}|E_{1}\rangle\otimes|E_{2}\rangle.$ (S28)
Note that the former one does not include components of $|E_{1}\rangle\otimes|E_{2}\rangle=|\lambda^{(0)}_{(1,1)}\rangle\!\rangle$. More generally, we find that the pure imaginary Liouvillean eigenmodes in layer $l$ have zero overlap with those $|\lambda^{(0)}_{(l,s)}\rangle\!\rangle$ where $s$ is odd.
Furthermore, we find that the LPTS breaking phenomenon exists only in layers where $(N-|l|)$ is even, i.e., where $l=l^{\prime\prime}$. Thus, studying the Liouvillean eigenmodes $\Big{\{}|\lambda^{(0)}_{(l,s)}\rangle\!\rangle\Big{\}}$ in layer $l^{\prime\prime}$ can help us to understand the nature of LPTS breaking. We define a special Liouvillean vector
$\big{|}\zeta\big{\rangle}\!\big{\rangle}=\sum_{l^{\prime\prime}}\frac{1}{\sqrt{N-1}}\Big{|}\lambda^{(0)}_{(l^{\prime\prime},1)}\Big{\rangle}\!\!\Big{\rangle},$ (S29)
where the $N-1$ normalizing factor counts the total number of layers with LPTS breaking. We confirmed, for the SCL model and $\gamma<\gamma_{\star}^{\mathrm{S}}$, the overlap between $\big{|}\zeta\big{\rangle}\!\big{\rangle}$ and the eigenmodes in the layer $l^{\prime\prime}$ is nearly zero. Once $\gamma>\gamma_{\star}^{\mathrm{S}}$, the overlap rises sharply to a finite number. This behavior mimics that of the real part of Liouvillean eigenvalues.
@@
The explicit form of $\mathcal{T}_{-}$ for the SCL model is
@@
where $\mathcal{K}$ is the Liouvillean complex-conjugation operator, $\mathcal{K}\mathcal{L}\mathcal{K}^{-1}=\mathcal{L}^{*}$, and $\hat{I}$ is the identity matrix. The factor $p_{x}=\pm 1$, which makes $\mathcal{T}_{-}^{2}=\pm 1$, respectively.
When $\gamma=0$, all eigenmodes are symmetric: $\lambda^{\prime}=-(\lambda^{\prime})^{\ast}$ and $\mathbb{P}\mathbb{T}|\lambda^{\prime}\rangle\!\rangle=|-(\lambda^{\prime})^{\ast}\rangle\!\rangle$. Thus, $|-(\lambda^{\prime})^{\star}\rangle\!\rangle$ is degenerate with $|\lambda^{\prime}\rangle\!\rangle$ and these lie on the imaginary axis. As the dephasing rate $\gamma$ increases until it exceeds the critical point $\gamma_{\star}^{\mathrm{S}}$, the symmetry of $\lambda^{\prime}=-(\lambda^{\prime})^{\ast}$ is broken, while the eigenvectors continue to be related via $\mathbb{P}\mathbb{T}|\lambda^{\prime}\rangle\!\rangle=|-(\lambda^{\prime})^{\star}\rangle\!\rangle$ remains. This indicates that, after the critical point, a few pairs of degenerate eigenvalues move symmetrically away from the imaginary axis along the real axis. In addition, a few eigenvalues stay at the imaginary axis regardless of the magnitude of $\gamma$. For each momentum subspace, these eigenmodes can be found using the condition of
@@
We note that, for the PXP model, we cannot find a $\mathcal{T}_{-}$ operator that satisfies Eq. (S30). Without $\mathcal{T}_{-}$ symmetry, the Liouvillean spectrum is intrinsically asymmetric about the imaginary axis [88], as confirmed by the numerical simulations.
Figure S2: The Liouvillean spectra of the effective non-Hermitian Hamiltonians for the PXP model and SCL model ($J_{\textrm{x},j}=0.1$) at two dephasing rates $\gamma=0.0001,0.01$. The black and blue dots indicate the spectra of thermal states, while the red dots denote the approximate scars. The system size is $L=12$. Other parameters are the same as the main text
Figure S3: (a)-(b) The Liouvillean spectra (left panel) and the overlap with $|\mathbb{Z}_{2}\rangle\otimes|\mathbb{Z}_{2}\rangle$ (right panel) corresponding to two dephasing rates, $\gamma=0.0004,0.04$, of the PXP model and DPXP model. The scar (S) and thermal (T) states are labeled according to $q=0,\pi$ momenta under translation, as indicated in the legend. (c)-(d) The case of the SCL model with $J_{\mathrm{x},j}=0,0.1$. The states are labeled by $p=\pm 1$, indicating the inversion symmetry is even (odd). The overlap is with the state $|\Pi\rangle\otimes|\Pi\rangle$, defined in the main text. In all plots, the black boundary dots indicate scar states with the local maximum overlap with the chosen initial state. The system size is $L=10$ in all plots.
Appendix S5 C. The origin of LPTS breaking
@@
In Fig. S2 we numerically argue that LPTS breaking due to dephasing does not exist in the non-Hermitian approximation which adds the term $i\gamma\sum^{L}_{j=1}\hat{\sigma}^{z}_{j}$ to the Hamiltonian. From the form of this term, it is clear that for the SCL model in the sector with conserved magnetization $\sum^{L}_{j=1}\hat{\sigma}^{z}_{j}=0$, the non-Hermitian operator in the spin $z$-direction has no effect on the Liouville spectrum. This is confirmed in Fig. S2, which also shows that the non-Hermitian PXP model immediately breaks LPTS by a very small $\gamma=10^{-4}$, in stark contrast with the results for its Liouvillean spectrum in the main text.
Appendix S6 D. Exact vs. approximate scars in SCL and PXP models
@@
For the SCL model, in the main text we emphasized a significant difference in the Liouvillean spectra of the two subspaces with reflection symmetry $p=\pm 1$, and also between the cases of approximate ($J_{\mathrm{x},j}\neq 0$) vs. exact ($J_{\mathrm{x},j}=0$) QMBSs. We highlight these differences in more detail in Fig. S3. In the odd subspace $p=-1$, most approximate scars exhibit LPTS under weak dissipation, while in the even subspace $p=+1$ they have immediately broken LPTS. Under strong dissipation, the behavior of the two sectors tends to be consistent. In addition, we also observe that the maximum overlaps of scars are uniformly distributed in the Liouvillean spectral range across different subspaces. In the case of weak dissipation, more scars in $p=-1$ odd subspace approach the zero point, $\mathrm{Re}(\lambda^{\prime}/\gamma)=0$. However, as the dephasing rate increases, they gradually move away from zero.
@@
For the PXP model, also shown in Fig. S3, we find qualitatively similar behavior to the SCL model, with a few minor differences. First, the LPTS breaking occurs in different momentum $q$ sectors under translation, rather than inversion symmetry $p$ in the SCL model. Second, there is no way to directly tune between approximate and exact scar limits. Instead, as a proxy for the latter, we consider a deformed PXP model (DPXP), introduced in the main text. With these provisos, the behavior of the PXP model largely mirrors that of SCL model, with the main difference that $\mathbb{PT}$ symmetry is only approximately obeyed, e.g., see the scar states with momentum $q=0$ in Fig. S3(a). As we explained above, this is attributed to a lack of $\mathcal{T}_{-}$ symmetry of the PXP Liouvillean. Nevertheles, these scarred eigenvalues still exhibit clustering around the imaginary axis at weak dephasing rates, giving rise to a certain level of robustness of QMBS signatures in dissipative dynamics, just like for the SCL model with $J_{\mathrm{x},j}\neq 0$.
Appendix S7 E. Further numerical results on Liouvillean spectra of different models
@@
Figure S4: 1D tilted Fermi-Hubbard (TFH) model, Eq. (S33), for two different dephasing rates. The gray dots indicate the Liouvillean spectra of thermal states, while the red dots indicate the spectra of scar states. The system size is $L=6$.
@@
Spontaneous LPTS breaking can also be observed in the approximate scar states of a 1D tilted Fermi-Hubbard (TFH) model, Fig. S4, whose Hamiltonian (after the Jordan-Wigner transformation) is [35]
@@
Figure S5: Hilbert space fragmentation in the frustrated Heisenberg ladder model (S34). (a) and (b) correspond to the case of $\eta=0$ and $\eta=0.01$, respectively. The gray dots indicate the Liouville spectra of all states, while the red dots denote the spectra of a few typical local HSF states. (c) The variation of the Liouvillean spectrum $\mathrm{Re}(\lambda^{\prime}/\gamma)$ of a few typical local HSF states with the dephasing rate $\gamma$. The red (bule) dots are $\eta=0.01$ (00). The system size is $L=8$ and the parameter $J=1$.

### v3

@@ Abstract
S1.4 A4. Difference between scar subspaces with and without a spectral transition
[…]
For the case $s\neq N-|l|-s$, degenerate perturbation theory is applicable when dissipation is very small. The general expression for the eigenvalue $\lambda_{(l,s)}$ at arbitrary order is given by sum-over-paths expansion:
$\lambda_{(l,s)}=i(2lJ+\Sigma_{(l,s)})+\sum_{n=1}^{\infty}\gamma_{\text{eff}}^{n}\sum_{\text{path}}\frac{\langle\lambda_{(l,s)}^{\mathrm{DF}}|\mathcal{D}_{l}^{\prime}|\lambda_{(l,p_{1})}^{\mathrm{DF}}\rangle\langle\lambda_{(l,p_{1})}^{\mathrm{DF}}|\mathcal{D}_{l}^{\prime}|\lambda_{(l,p_{2})}^{\mathrm{DF}}\rangle\cdots\langle\lambda_{(l,p_{n-1})}^{\mathrm{DF}}|\mathcal{D}_{l}^{\prime}|\lambda_{(l,N-|l|-s)}^{\mathrm{DF}}\rangle}{\prod_{j=1}^{n-1}(i\Sigma_{(l,s)}-i\Sigma_{(l,p_{j})})},$ (S45)
where $\left|p_{j}\right\rangle$ are the intermediate states between $\left|\lambda_{(l,s)}^{\mathrm{DF}}\right\rangle$ and $\left|\lambda_{(l,N-|l|-s)}^{\mathrm{DF}}\right\rangle$.
The nature of the correction term depends critically on the order $n$. The shortest path between $\left|\lambda_{(l,s)}^{\mathrm{DF}}\right\rangle$ and $\left|\lambda_{(l,N-|l|-s)}^{\mathrm{DF}}\right\rangle$ costs |N|l|2s||N-|l|-2s| steps, which means only odd (even) order terms make a contribution when N|l|N-|l| is odd (even). Consequently, the eigenvalues are all purely imaginary when N|l|N-|l| is even, while they all acquire nonzero real part when N|l|N-|l| is odd. For example, in the case $N=5,l=+3,s=0$, only the second-order term should be considered:
$\gamma_{\text{eff}}^{2}\frac{\langle\lambda_{(3,0)}^{\mathrm{DF}}|\mathcal{D}_{l}^{\prime}|\lambda_{(3,1)}^{\mathrm{DF}}\rangle\langle\lambda_{(3,1)}^{\mathrm{DF}}|\mathcal{D}_{l}^{\prime}|\lambda_{(3,2)}^{\mathrm{DF}}\rangle}{(i\Sigma_{(3,0)}-i\Sigma_{(3,1)})}\in i\mathbb{R},$ (S46)
which is purely imaginary. While in the case $N=5,l=+2,s=0$, the third-order term dominates:
$\gamma_{\text{eff}}^{3}\frac{\langle\lambda_{(2,0)}^{\mathrm{DF}}|\mathcal{D}_{l}^{\prime}|\lambda_{(2,1)}^{\mathrm{DF}}\rangle\langle\lambda_{(2,1)}^{\mathrm{DF}}|\mathcal{D}_{l}^{\prime}|\lambda_{(2,2)}^{\mathrm{DF}}\rangle\langle\lambda_{(2,2)}^{\mathrm{DF}}|\mathcal{D}_{l}^{\prime}|\lambda_{(2,3)}^{\mathrm{DF}}\rangle}{(i\Sigma_{(2,0)}-i\Sigma_{(2,1)})(i\Sigma_{(2,0)}-i\Sigma_{(2,2)})}\in\mathbb{R},$ (S47)
which is real.
For the case s=N|l|ss=N-|l|-s, we should consider the non-degenerate expansion. Here, any path connecting the state back to itself must involve an even number of steps under the dissipative operator $\mathcal{D}_{l}^{\prime}$, consequently forcing all odd-order corrections to vanish identically. Its eigenvalue correction is given by:
$\lambda_{(l,s)}=i\Sigma_{(l,s)}+\sum_{n=1}^{\infty}\gamma_{\text{eff}}^{2n}\sum_{\text{path}}\frac{\langle\lambda_{(l,s)}^{\mathrm{DF}}|\mathcal{D}_{l}^{\prime}|\lambda_{(l,p_{1})}^{\mathrm{DF}}\rangle\langle\lambda_{(l,p_{1})}^{\mathrm{DF}}|\mathcal{D}_{l}^{\prime}|\lambda_{(l,p_{2})}^{\mathrm{DF}}\rangle\cdots\langle\lambda_{(l,p_{2n-1})}^{\mathrm{DF}}|\mathcal{D}_{l}^{\prime}|\lambda_{(l,s)}^{\mathrm{DF}}\rangle}{\prod_{j=1}^{2n-1}(i\Sigma_{(l,s)}-i\Sigma_{(l,p_{j})})}.$ (S48)
Noticing the factor $i^{2n-1}$ in the denominator, the eigenvalue should be purely imaginary when dissipation is small.
The key distinction between two subspaces emerges from the parity of coupling path length. Within the $\mathrm{S},{-}1$’ subspace, coupling between degenerate states necessitates an even number of applications of the dissipative operator $\mathcal{D}_{l}^{\prime}$. Consequently, for small $\gamma$, only even-order corrections are non-zero, resulting in the eigenvalue shifting purely along the imaginary axis. Conversely, in the $\mathrm{S},{+}1$’ subspace, the coupling mechanism allows odd-order corrections, leading to a shift along the real axis even for small dissipation. This distinction leads to the fact that the eigenvalues in $\mathrm{S},{-}1$’ subspace undergoes a spectral transition while the eigenvalues in $\mathrm{S},{+}1$’ do not.
To conclude, both the thermal subspace and the bath contributes to the eigenvalue correction in $\mathrm{S},{\pm}1$’ subspace. However, only the eigenvalues in $\mathrm{S},{-}1$’ subspace undergo spectral transition due to the parity of coupling path length. Additionally, in the PXP model, the same analysis can be used to prove that the spectral transition only occurs exclusively in $\mathrm{S},0$’ subspace.
S2 B. Symmetries of the Liouvillian
First, we briefly analyze the symmetries of the CL model and the PXP model in an open quantum setting. We denote the basis states for the Liouvillian $\mathcal{L}$ of the CL model by
$|B\rangle\!\rangle=\left|\begin{subarray}{c}u_{1}\\ d_{1}\end{subarray}...\begin{subarray}{c}u_{i}\\ d_{i}\end{subarray}...\begin{subarray}{c}u_{N}\\ d_{N}\end{subarray}\right\rangle\!\otimes\!|\begin{subarray}{c}u^{\prime}_{1}\\ d^{\prime}_{1}\end{subarray}...\begin{subarray}{c}u^{\prime}_{i}\\ d^{\prime}_{i}\end{subarray}...\begin{subarray}{c}u^{\prime}_{N}\\ d^{\prime}_{N}\end{subarray}\rangle,\quad u,d,u^{\prime},d^{\prime}\in\{\uparrow,\downarrow\}.$ (S49)
The inversion symmetry superoperator $\mathcal{P}$ has the following action on basis states:
$\mathcal{P}|B\rangle\!\rangle=|\bar{B}\rangle\!\rangle=\left|\begin{subarray}{c}\bar{u}_{1}\\ \bar{d}_{1}\end{subarray}...\begin{subarray}{c}\bar{u}_{i}\\ \bar{d}_{i}\end{subarray}...\begin{subarray}{c}\bar{u}_{N}\\ \bar{d}_{N}\end{subarray}\right\rangle\!\otimes\!|\begin{subarray}{c}\bar{u}^{\prime}_{1}\\ \bar{d}^{\prime}_{1}\end{subarray}...\begin{subarray}{c}\bar{u}^{\prime}_{i}\\ \bar{d}^{\prime}_{i}\end{subarray}...\begin{subarray}{c}\bar{u}^{\prime}_{N}\\ \bar{d}^{\prime}_{N}\end{subarray}\rangle,$ (S50)
where $\bar{u}_{i}$ ($\bar{d}_{i}$) denotes the $i$th spin state opposite to $u_{i}$ ($d_{i}$). We denote the eigenvalues of $\mathcal{P}$ by $p=\pm 1$.
In addition, we should consider the reflection superoperator $\mathcal{R}$,
$\mathcal{R}|B\rangle\!\rangle=|B^{\prime}\rangle\!\rangle=\left|\begin{subarray}{c}u_{N}\\ d_{N}\end{subarray}...\begin{subarray}{c}u_{i}\\ d_{i}\end{subarray}...\begin{subarray}{c}u_{1}\\ d_{1}\end{subarray}\right\rangle\!\otimes\!|\begin{subarray}{c}u^{\prime}_{N}\\ d^{\prime}_{N}\end{subarray}...\begin{subarray}{c}u^{\prime}_{i}\\ d^{\prime}_{i}\end{subarray}...\begin{subarray}{c}u^{\prime}_{1}\\ d^{\prime}_{1}\end{subarray}\rangle.$ (S51)
The CL Hamiltonian of the two horizontal legs $H_{\rm{h},\alpha}$ remains unchanged under the action of inversion symmetry operator $\mathcal{P}$, but it maps to $H_{\rm{h},\alpha}$ under the action of reflection symmetry operator $\mathcal{R}$ when the number of spins $N$ is odd. Hence, we will restrict to the cases when $N$ is even, when the eigenmodes can be classified according to the eigenvalues $r=\pm 1$ of $\mathcal{R}$.
On the other hands, for the PXP model, we denote the basis states for the Liouvillian $\mathcal{L}$ by $|B\rangle\!\rangle=\left|u_{1}...u_{i}...u_{N}\right\rangle\!\otimes\!|u^{\prime}_{1}...u^{\prime}_{i}...u^{\prime}_{N}\rangle$. The translation symmetry superoperator $\mathcal{Q}$ is defined by
$\mathcal{Q}|B\rangle\!\rangle=|B^{\prime}\rangle\!\rangle=\left|u_{N}...u_{i+1}...u_{N-1}\right\rangle\!\otimes\!|u^{\prime}_{N}...u^{\prime}_{i+1}...u^{\prime}_{N-1}\rangle,$ (S52)
which we use to characterize the eigenmodes in terms of conserved momentum $q$.
@@
Next, we study the symmetry properties of the Liouvillian eigenmodes responsible for the spectral transition discussed in the main text. Recall that the unperturbed $\mathcal{S}_{0}$ has eigenmodes $|\lambda_{(l,s)}^{(0)}\rangle\!\rangle=|E_{l+s}^{\mathrm{S}}\rangle\otimes|E_{s}^{\mathrm{S}}\rangle$. Take $N=3$ case as an example. There are $N+1=4$ eigenstates of $\hat{S}_{0}$ in the maximal spin representation:
$|E_{0}^{\mathrm{S}}\rangle=|---\rangle,\quad\quad|E_{1}^{\mathrm{S}}\rangle=\frac{1}{\sqrt{3}}\Big(|+--\rangle+|-+-\rangle+|--+\rangle\Big),$ (S53)
$|E_{2}^{\mathrm{S}}\rangle=\frac{1}{\sqrt{3}}\Big(|++-\rangle+|+-+\rangle+|-++\rangle\Big),\quad\quad|E_{3}^{\mathrm{S}}\rangle=|+++\rangle,$
@@
with energies 3J-3J, J-J, $J$ and $3J$, respectively. The Liouvillian spectrum of $\mathcal{S}_{0}$ has a seven-layer structure with (imaginary) eigenvalues 0, $\pm 2J$, $\pm 4J$, $\pm 6J$.
@@
The Liouvillian spectrum of $(\mathcal{S}_{0}+\gamma\mathcal{D}^{\prime})$ maintains the seven-layer structure, but now includes eigenvalues with non-zero real parts. In the layer of $l=1$, there are two types of Liouvillian eigenmodes. Both of them are the linear superpositions of $|E_{0}^{\mathrm{S}}\rangle\otimes|E_{1}^{\mathrm{S}}\rangle$, $|E_{1}^{\mathrm{S}}\rangle\otimes|E_{2}^{\mathrm{S}}\rangle$, and $|E_{2}^{\mathrm{S}}\rangle\otimes|E_{3}^{\mathrm{S}}\rangle$, but with different coefficients. One of them is purely imaginary,
$\frac{1}{\sqrt{2}}|E_{0}^{\mathrm{S}}\rangle\otimes|E_{1}^{\mathrm{S}}\rangle-\frac{1}{\sqrt{2}}|E_{2}^{\mathrm{S}}\rangle\otimes|E_{3}^{\mathrm{S}}\rangle,$ (S54)
@@
$\frac{1}{2}|E_{0}^{\mathrm{S}}\rangle\otimes|E_{1}^{\mathrm{S}}\rangle+\frac{1}{2}|E_{2}^{\mathrm{S}}\rangle\otimes|E_{3}^{\mathrm{S}}\rangle\pm\frac{1}{\sqrt{2}}|E_{1}^{\mathrm{S}}\rangle\otimes|E_{2}^{\mathrm{S}}\rangle.$ (S55)
Note that the former one does not include components of $|E_{1}^{\mathrm{S}}\rangle\otimes|E_{2}^{\mathrm{S}}\rangle=|\lambda^{(0)}_{(1,1)}\rangle\!\rangle$. More generally, we find that the pure imaginary Liouvillian eigenmodes in layer $l$ have zero overlap with those $|\lambda^{(0)}_{(l,s)}\rangle\!\rangle$ where $s$ is odd.
Furthermore, we find that the spectral transition phenomenon exists only in layers where (N|l|)(N-|l|) is even, i.e., where $l=-N,-N+2,\ldots,N$. Thus, studying the Liouvillian eigenmodes $\Big\{|\lambda^{(0)}_{(l,s)}\rangle\!\rangle\Big\}$ in such layers can help us to understand the nature of Liouvillian spectral transition. We define a special Liouvillian vector
$\big|\zeta\big\rangle\!\big\rangle=\sum_{(N-|l|)\in\mathrm{even}}\frac{1}{\sqrt{N-1}}\Big|\lambda^{(0)}_{(l,1)}\Big\rangle\!\!\Big\rangle,$ (S56)
where the N1N-1 normalizing factor counts the total number of layers with spectral transition. We confirmed, for the CL model and $\gamma<\gamma_{\star}^{\mathrm{S}}$, the overlap between $\big|\zeta\big\rangle\!\big\rangle$ and the eigenmodes in the layer $l^{\prime}$ is nearly zero. Once $\gamma>\gamma_{\star}^{\mathrm{S}}$, the overlap rises sharply to a finite number. This behavior mimics that of the real part of Liouvillian eigenvalues.
@@
The explicit form of $\mathcal{T}_{-}$ for the CL model is
@@
where $\mathcal{K}$ is the Liouvillian complex-conjugation operator, $\mathcal{K}\mathcal{L}\mathcal{K}^{-1}=\mathcal{L}^{*}$, and $\hat{I}$ is the identity matrix. The factor $p_{x}=\pm 1$, which makes $\mathcal{T}_{-}^{2}=\pm 1$, respectively.
When $\gamma=0$, $|-(\lambda^{\prime})^{\ast}\rangle\!\rangle$ and $|\lambda^{\prime}\rangle\!\rangle$ represent the same state which lies on the imaginary axis. As the dephasing rate $\gamma$ increases until it exceeds the critical point $\gamma_{\star}^{\mathrm{S}}$, the symmetry of $\lambda^{\prime}=-(\lambda^{\prime})^{\ast}$ is broken, while the eigenvectors continue to be related via $\mathbb{P}\mathbb{T}|\lambda^{\prime}\rangle\!\rangle=|-(\lambda^{\prime})^{\star}\rangle\!\rangle$ remains. This indicates that, after the critical point, a few pairs of degenerate eigenvalues move symmetrically away from the imaginary axis along the real axis. In addition, a few eigenvalues stay at the imaginary axis regardless of the magnitude of $\gamma$. These eigenmodes can be found using the condition of
@@
We note that, for the PXP model, we cannot find a $\mathcal{T}_{-}$ operator that satisfies Eq. (S57). Without $\mathcal{T}_{-}$ symmetry, the Liouvillian spectrum is intrinsically asymmetric about the imaginary axis [89], as confirmed by the numerical simulations.
Figure S3: The Liouvillian spectra of the effective non-Hermitian Hamiltonians for the PXP model and CL model ($J_{\textrm{x},j}=0.1$) at two dephasing rates $\gamma=0.0001,0.01$. The black and blue dots indicate the spectra of thermal states, while the red dots denote the approximate scars. The system size is $L=12$. Other parameters are the same as the main text
Figure S4: (a)-(b) The Liouvillian spectra (left panel) and the overlap with $|\mathbb{Z}_{2}\rangle\otimes|\mathbb{Z}_{2}\rangle$ (right panel) corresponding to two dephasing rates, $\gamma=0.0004,0.04$, of the PXP model and DPXP model. The scar (S) and thermal (T) states are labeled according to $q=0,\pi$ momenta under translation, as indicated in the legend. (c)-(d) The case of the CL model with $J_{\mathrm{x},j}=0,0.1$. The states are labeled by $p=\pm 1$, indicating the inversion symmetry is even (odd). The overlap is with the state $|\Pi\rangle\otimes|\Pi\rangle$, defined in the main text. In all plots, the black boundary dots indicate scar states with the local maximum overlap with the chosen initial state. The system size is $L=10$ in all plots.
S3 C. The origin of Liouvillian spectrum transition
@@
In Fig. S3 we numerically argue that spectrum transition due to dephasing does not exist in the non-Hermitian approximation which adds the term $i\gamma\sum^{L}_{j=1}\hat{\sigma}^{z}_{j}$ to the Hamiltonian. From the form of this term, it is clear that for the CL model in the sector with conserved magnetization $\sum^{L}_{j=1}\hat{\sigma}^{z}_{j}=0$, the non-Hermitian operator in the spin $z$-direction has no effect on the Liouville spectrum. This is confirmed in Fig. S3, which also shows that the non-Hermitian PXP model immediately spectrum transition by a very small $\gamma=10^{-4}$, in stark contrast with the results for its Liouvillian spectrum in the main text.
S4 D. Exact vs. approximate scars in CL and PXP models
@@
For the CL model, in the main text we emphasized a significant difference in the Liouvillian spectra of the two inversion symmetry subspaces $p=\pm 1$, and also between the cases of approximate ($J_{\mathrm{x},j}\neq 0$) vs. exact ($J_{\mathrm{x},j}=0$) QMBSs. We highlight these differences in more detail in Fig. S4(c)-(d). In the odd subspace p=1p=-1, most approximate scars exhibit spectrum transition under weak dissipation, while in the even subspace $p=+1$ they have immediately spectrum transition. Under strong dissipation, the behavior of the two sectors tends to be consistent. In addition, we also observe that the maximum overlaps of scars are uniformly distributed in the Liouvillian spectral range across different subspaces. In the case of weak dissipation, more scars in p=1p=-1 odd subspace approach the zero point, $\mathrm{Re}(\lambda^{\prime}/\gamma)=0$. However, as the dephasing rate increases, they gradually move away from zero.
@@
For the PXP model, also shown in Fig. S4(a)-(b), we find qualitatively similar behavior to the CL model, with a few minor differences. First, the spectrum transition occurs in different momentum $q$ sectors under translation, rather than inversion symmetry $p$ in the CL model. Second, there is no way to directly tune between approximate and exact scar limits. Instead, as a proxy for the latter, we consider a deformed PXP model (DPXP), introduced in the main text. With these provisos, the behavior of the PXP model largely mirrors that of CL model, with the main difference that $\mathbb{PT}$ symmetry is only approximately obeyed, e.g., see the scar states with momentum $q=0$ in Fig. S4(a). As we explained above, this is attributed to a lack of $\mathcal{T}_{-}$ symmetry of the PXP Liouvillian. Nevertheles, these scarred eigenvalues still exhibit clustering around the imaginary axis at weak dephasing rates, giving rise to a certain level of robustness of QMBS signatures in dissipative dynamics, just like for the CL model with $J_{\mathrm{x},j}\neq 0$.
S5 E. Nsumerical results on Liouvillian spectra of different models
@@
Figure S5: 1D tilted Fermi-Hubbard (TFH) model, Eq. (S60), for two different dephasing rates. The gray dots indicate the Liouvillian spectra of thermal states, while the red dots indicate the spectra of scar states. The system size is $L=6$.
@@
Liouvillian spectrum transition can also be observed in the approximate scar states of a 1D tilted Fermi-Hubbard (TFH) model, Fig. S5, whose Hamiltonian (after the Jordan-Wigner transformation) is [35]
@@
Figure S6: Hilbert space fragmentation in the frustrated Heisenberg ladder model (S61). (a) and (b) correspond to the case of $\eta=0$ and $\eta=0.01$, respectively. The gray dots indicate the Liouville spectra of all states, while the red dots denote the spectra of a few typical local HSF states. (c) The variation of the Liouvillian spectrum $\mathrm{Re}(\lambda^{\prime}/\gamma)$ of a few typical local HSF states with the dephasing rate $\gamma$. The red (bule) dots are $\eta=0.01$ (0). The system size is $L=8$ and the parameter $J=1$.